### v2

# MSTN: Fast and Efficient Multivariate Time Series Prediction Model

Abstract
Real-world time series often exhibit strong non-stationarity, complex nonlinear dynamics, and behaviour expressed across multiple temporal scales, from rapid local fluctuations to slow-evolving long-range trends. However, many contemporary architectures impose rigid, fixed-scale structural priors—such as patch-based tokenization, predefined receptive fields, or frozen backbone encoders—which can over-regularize temporal dynamics and limit adaptability to abrupt high-magnitude events. To handle this, we introduce the Multi-scale Temporal Network (MSTN), a hybrid neural architecture grounded in an Early Temporal Aggregation principle. MSTN integrates three complementary components: (i) a multi-scale convolutional encoder that captures fine-grained local structure; (ii) a sequence modeling module that learns long-range dependencies through either recurrent or attention-based mechanisms; and (iii) a self-gated fusion stage incorporating squeeze–excitation and multi-head attention to dynamically modulate cross-scale representations. This design enables MSTN to flexibly model temporal patterns spanning milliseconds to extended horizons, while avoiding the computational burden typically associated with long-context models. Across extensive benchmarks covering forecasting, imputation, classification, and cross-dataset generalization, MSTN consistently delivers state-of-the-art performance, outperforming recent leading approaches including TIME-LLM, HiMTM, SOFTS, LLM4TS, TimesNet, and PatchTST, and establishing new best results on 24 out of 32 datasets. Despite its strong performance, MSTN remains lightweight and supports fast inference, making it well suited for deployment on edge devices and resource-constrained environments.
keywords:
Multivariate time series , multi-scale temporal modeling , time series forecasting , classification , imputation , computational efficiency , deep learning architectures
[…]
[inst1]organization=Department of Computer Science and Engineering, Indian Institute of Technology Indore, addressline=Indore, postcode=452020, state=Madhya Pradesh, country=India
1 Introduction
Multivariate time series (MTS) analysis has been a classical problem with significant real-world applications including healthcare monitoring [10.1145/3531326], traffic management [kadiyala2014multivariate], environment [gruca2022weather4cast], industrial predictive maintenance, energy  [6714975] and behavioral analytics. It has been extensively studied for many years and has been widely used in tasks such as weather forecasting, imputing missing data for data mining, and classification of trajectories for action recognition [wu2022autoformer, zhao2024himtmhierarchicalmultiscalemasked]. Due to its significant practical value, time series analysis has attracted considerable attention. Unlike language or video, time-series data consists of individual scalar measurements, where single time points rarely carry sufficient semantic meaning. This fundamental characteristic forces models to extract information from temporal variation patterns of continuity, periodicity, and trend, making robust temporal modeling exceptionally challenging [wu2023timesnettemporal2dvariationmodeling, lim2021time].
Deep learning (DL) models have demonstrated strong performance not only in forecasting tasks, but also in representation learning, where abstract features can be extracted and transferred to downstream tasks such as classification and forecasting, achieving state-of-the-art (SOTA) results [nie2023timeseriesworth64]. Recent findings show that simple linear models can outperform sophisticated Transformer variants on common benchmarks [zeng2022transformerseffectivetimeseries], highlighting fundamental limitations in contemporary architectures. The design of temporal models faces the constant challenge of balancing model strength with computational efficiency. Models such as convolutional neural networks (CNNs) [franceschi2019unsupervised, bai2018empirical] and recurrent neural networks (RNNs) with their gated variants [6795963, lai2018modelinglongshorttermtemporal] suffer from vanishing gradients and limited parallelism, which has shifted the paradigm towards more complex models. Temporal convolutional networks (TCN)  [he2019temporal] improve computational efficiency but are constrained by fixed receptive fields. The Transformer architecture  [vaswani2023attentionneed] originally promised global dependency modeling, but its quadratic complexity prompted a line of research focused primarily on sparsification. Models like Informer [zhou2021informer] and FEDformer [zhou2022fedformer] develop efficient attention variants, yet often retain a point-wise view of time series, treating each time step as an independent token and struggling to capture local semantic structures.
These challenges have shaped several predominant, yet ultimately constrained, modern paradigms. These can be formalized into three primary Channel Strategies (CS) [qiu2025comprehensivesurveydeeplearning]: Channel Dependence (CD), which favors high feature completeness at a computational cost; Channel Independence (CI), which prioritizes speed by isolating channels; and Channel Partiality (CP), which aims for a balance by reducing channel complexity. CI based models like PatchTST [nie2023timeseriesworth64] improves local semantic modeling through patch tokenization but introduces rigidity through fixed patch sizes; LLM-derived approaches such as LLM4TS [10.1145/3719207] and TIME-LLM [jin2024timellmtimeseriesforecasting] depend on frozen backbone architectures. CD based models such as TimesNet [wu2023timesnettemporal2dvariationmodeling] transforms time series into periodic 2D structures but depends heavily on accurate period discovery, limiting robustness for irregular or non-stationary signals. MTST [zhang2024multiresolutiontimeseriestransformerlongterm] models multiple resolutions but relies on fixed resolution sets and may dismiss fine-grained dynamics. Even highly efficient recent systems such as SOFTS [han2024softsefficientmultivariatetime] prioritize channel efficiency through linear factorization, inherently sacrificing temporal completeness and weakening multi-scale temporal representation. Similarly, the CP Channels strategy in MCformer [10533212] achieves efficiency by partially blending a specific number of channels, but this partiality limits the recovery of full inter-channel context. Consequently, such architectures struggle to sustain fidelity in complex long-sequence environments.
We identify a fundamental limitation on these approaches: the imposition of single-scale architectural assumptions such as fixed patch sizes, fixed periodicity-driven 2D mappings, struggle to capture discriminative cues manifesting across diverse temporal horizons, or static backbone structures. However, real-world temporal phenomena are inherently multi-scale: micro-events, meso-scale fluctuations, and macro trends often coexist. Consequently, existing models struggle to capture discriminative cues manifesting across diverse temporal horizons, particularly in domains such as risky driving behaviour, physiological safety monitoring, or mechanical prognostics—where multi-scale variability is intrinsic. This gap motivates a more flexible, hierarchy-aware temporal representation without relying on a single predefined structure.
To address these limitations, we propose the Multi-scale Temporal Network (MSTN), a hybrid architecture founded on the Early Temporal Aggregation (ETA) principle. MSTN naturally captures hierarchical temporal patterns without fixed structural assumptions through its central innovation: ETA condenses the full temporal input into a single static feature vector via an $L\rightarrow 1$ transformation, effectively neutralizing the $\mathcal{O}(L^{2})$ inference bottleneck while preserving global completeness. This design allows the dual encoder to retain $\mathcal{O}(L^{2})$ modeling capacity during representation extraction, while all subsequent refinement operates in guaranteed constant-time $\mathcal{O}(1)$ complexity. MSTN unifies multi-scale feature extraction with global temporal reasoning through a dual-path architecture: convolutional encoders capture fine-grained local structure, while a Bidirectional Long Short-Term Memory (BiLSTM) or Transformer backbone models long-range dependencies. Their outputs are fused through a self-gated fusion (SGF) mechanism enriched by squeeze-and-excitatio (SE) recalibration and multi-head attention (MHA), producing compact, discriminative, and temporally coherent representations. We empirically validate the MSTN framework through two realizations: MSTN–BiLSTM, ideal for smooth continuous behaviors, and MSTN–Transformer, suitable for complex, long-range event modeling.
We demonstrate that MSTN establishes new SOTA performance on 24 of 32 benchmark datasets for time series analysis. This includes 9 forecasting benchmarks, 6 imputation tasks, and 17 classification datasets (10 from the standard UEA archive and 7 additional international benchmarks), demonstrating its versatility and improved performance across diverse applications. MSTN establishes a new SOTA in long-term forecasting, attaining the lowest prediction error on 6 out of 9 standard benchmarks and consistently exceeds leading architectures, including EMTSF [alharthi2025emtsfextraordinarymixturesotamodels], LLM4TS [10.1145/3719207], HiMTM [zhao2024himtmhierarchicalmultiscalemasked], TIME-LLM [jin2024timellmtimeseriesforecasting], MTST [zhang2024multiresolutiontimeseriestransformerlongterm], GPT2(6) [zhou2023fitsallpowergeneraltime], SOFTS [han2024softsefficientmultivariatetime] iTransformer [liu2024itransformerinvertedtransformerseffective], and PatchTST [nie2023timeseriesworth64]. Complementing these results, the MSTN-Transformer variant achieves first-place rankings 31 out of 48 ($\approx$65%) in all 48 imputation scenarios across the ETT, Electricity, and Weather datasets, demonstrating strong resilience under up to 50% missingness. In the classification task, MSTN achieves improved performance over recent SOTA architectures such as GTP2(6) [zhou2023fitsallpowergeneraltime], TimesNet [wu2023timesnettemporal2dvariationmodeling], LigthtTS  [zhang2022morefastmultivariatetime] and FEDformer [zhou2022fedformer]. We further evaluate MSTN for generalizability across 7 benchmark datasets from various domains like human safety, physical activity recognition, animal-welfare monitoring, and mechanical prognostics. Without domain-specific tuning, MSTN achieves improved performance, ranking first across all datasets, and illustrates its robustness under diverse temporal dynamics. Concretely, MSTN-Transformer achieves unprecedented performance with 2.87 MB memory footprint, and 0.72 ms inference latency, representing a $22.4\times$ accuracy improvement and $34.7\times$ speedup over SOFTS [han2024softsefficientmultivariatetime]. The model further demonstrates improved classification accuracy (99.53%) with ultra-low latency (0.15 ms), significantly outperforming TimesNet (99.18%, 5.07 ms) and PatchTST (98.64%, 4.80 ms) on the Rodegast et al. [DARUS33012023] dataset. While maintaining compact model sizes (0.54–7.14 MB), this balance of SOTA accuracy, robustness, and efficiency positions MSTN as a preferred solution for real-time and edge-AI temporal modeling applications.
2 Results
This section presents a comprehensive evaluation of the proposed MSTN model. We demonstrate its performance on three time-series tasks: (1) long-term forecasting, (2) imputation, (3) classification. Additionally, we also present the generalizability study on standard time-series benchmark datasets.
2.1 Benchmark Datasets
Table 1: Benchmark datasets and imputation performance comparison.
(a) Benchmark datasets for forecasting, imputation, and classification (10 UEA).
Task Datasets Features Timesteps Metrics Length
Imputation ETTh1/ETTh2 7 17,420 MSE, MAE 96
ETTm1/ETTm2 7 69,680
Electricity 321 26,304
Weather 21 52,696
Forecasting ETTh1/ETTh2 7 17,420 MSE, MAE 96–720
ETTm1/ETTm2 7 69,680
Electricity 321 26,304
Traffic 862 17,544
Weather 21 52,696
ILI 7 966
Classification EthanolConcentration 3 1,751 Accuracy 29-1,751
FaceDetection 144 62
Handwriting 3 152
Heartbeat 61 405
JapaneseVowels 12 29
PEMS-SF 963 144
SelfRegulationSCP1 6 896
SelfRegulationSCP2 7 1,152
SpokenArabicDigits 13 93
UWaveGestureLibrary 3 315
(b) Imputation task results. Random masking ratios of 12.5%, 25%, 37.5%, and 50% are applied. Red and blue denote first and second best results, respectively.
Models MSTN-Tra. MSTN BiL. GPT2(3) TimesNet PatchTST LightTS DLinear Stationary
(Ours) (Ours) (2023) (2023) (2023) (2022) (2023) (2022)
Mask R. MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
ETTh1
12.5% 0.024 0.040 0.032 0.047 0.043 0.140 0.057 0.159 0.093 0.201 0.240 0.345 0.151 0.267 0.060 0.165
25% 0.050 0.083 0.067 0.096 0.054 0.156 0.069 0.178 0.107 0.217 0.265 0.364 0.180 0.292 0.080 0.189
37.5% 0.083 0.131 0.103 0.147 0.072 0.180 0.084 0.196 0.120 0.230 0.296 0.382 0.215 0.318 0.102 0.212
50% 0.127 0.187 0.148 0.204 0.107 0.216 0.102 0.215 0.141 0.248 0.334 0.404 0.257 0.347 0.133 0.240
Avg 0.071 0.110 0.088 0.124 0.069 0.173 0.078 0.187 0.115 0.224 0.284 0.373 0.201 0.306 0.094 0.201
ETTh2
12.5% 0.009 0.024 0.022 0.040 0.039 0.125 0.040 0.130 0.057 0.152 0.101 0.231 0.100 0.216 0.042 0.133
25% 0.022 0.054 0.043 0.078 0.044 0.135 0.046 0.141 0.061 0.158 0.115 0.246 0.127 0.247 0.049 0.147
37.5% 0.038 0.089 0.069 0.121 0.051 0.147 0.052 0.151 0.067 0.166 0.126 0.257 0.158 0.276 0.056 0.158
50% 0.066 0.135 0.102 0.170 0.059 0.158 0.060 0.162 0.073 0.174 0.136 0.268 0.183 0.299 0.065 0.170
Avg 0.034 0.076 0.059 0.102 0.048 0.141 0.049 0.146 0.065 0.163 0.119 0.250 0.142 0.259 0.053 0.152
ETTm1
12.5% 0.024 0.038 0.036 0.049 0.017 0.085 0.019 0.092 0.041 0.130 0.075 0.180 0.058 0.162 0.026 0.107
25% 0.050 0.080 0.072 0.099 0.022 0.096 0.023 0.101 0.044 0.135 0.093 0.206 0.080 0.193 0.032 0.119
37.5% 0.082 0.124 0.112 0.150 0.029 0.111 0.029 0.111 0.049 0.143 0.113 0.231 0.103 0.219 0.039 0.131
50% 0.121 0.176 0.160 0.210 0.040 0.128 0.036 0.124 0.055 0.151 0.134 0.255 0.132 0.248 0.047 0.145
Avg 0.069 0.104 0.095 0.127 0.028 0.105 0.027 0.107 0.047 0.140 0.104 0.218 0.093 0.206 0.036 0.126
ETTm2
12.5% 0.039 0.048 0.052 0.057 0.017 0.076 0.018 0.080 0.026 0.094 0.034 0.127 0.062 0.166 0.021 0.088
25% 0.080 0.097 0.107 0.116 0.020 0.080 0.020 0.085 0.028 0.099 0.042 0.143 0.085 0.196 0.024 0.096
37.5% 0.132 0.153 0.164 0.176 0.022 0.087 0.023 0.091 0.030 0.104 0.051 0.159 0.106 0.222 0.027 0.103
50% 0.187 0.215 0.225 0.241 0.025 0.095 0.026 0.098 0.034 0.110 0.059 0.174 0.131 0.247 0.030 0.108
Avg 0.109 0.128 0.137 0.147 0.021 0.084 0.022 0.088 0.029 0.102 0.046 0.151 0.096 0.208 0.026 0.099
ECL
12.5% 0.002 0.014 0.034 0.048 0.080 0.194 0.085 0.202 0.055 0.160 0.102 0.229 0.092 0.214 0.093 0.210
25% 0.005 0.028 0.069 0.096 0.087 0.203 0.089 0.206 0.065 0.175 0.121 0.252 0.118 0.247 0.097 0.214
37.5% 0.008 0.044 0.108 0.147 0.094 0.211 0.094 0.213 0.076 0.189 0.141 0.273 0.144 0.276 0.102 0.220
50% 0.013 0.063 0.150 0.202 0.101 0.220 0.100 0.221 0.091 0.208 0.160 0.293 0.175 0.305 0.108 0.228
Avg 0.007 0.037 0.090 0.123 0.090 0.207 0.092 0.210 0.072 0.183 0.131 0.262 0.132 0.260 0.100 0.218
Weather
12.5% 0.003 0.016 0.026 0.036 0.026 0.049 0.025 0.045 0.029 0.049 0.047 0.101 0.039 0.084 0.027 0.051
25% 0.007 0.032 0.051 0.070 0.028 0.052 0.029 0.052 0.031 0.053 0.052 0.111 0.048 0.103 0.029 0.056
37.5% 0.011 0.049 0.078 0.109 0.033 0.060 0.031 0.057 0.035 0.058 0.058 0.121 0.057 0.117 0.033 0.062
50% 0.015 0.068 0.111 0.153 0.037 0.065 0.034 0.062 0.038 0.063 0.065 0.133 0.066 0.134 0.037 0.068
Avg 0.009 0.041 0.066 0.092 0.031 0.056 0.030 0.054 0.033 0.056 0.055 0.117 0.052 0.110 0.032 0.059
We evaluate the proposed method across three fundamental time series tasks: forecasting, missing value imputation, and classification using established public benchmark datasets. Table 1 provides a detailed overview of all datasets. Missing value imputation is performed under a Missing Completely at Random (MCAR) setting, with masking ratios ranging from 12.5%, 25%, 37.5% and 50%. Benchmarks include ETTh1, ETTh2, ETTm1, ETTm2, Electricity, and Weather datasets [zhou2021informer, electricityloaddiagrams20112014_321, Kolle2025_BGCJenaWeather]. Imputation results are quantified using Mean Squared Error (MSE) and Mean Absolute Error (MAE) over sequence lengths of 96 time steps.
Long-term forecasting performance is assessed on nine widely-used benchmarks: ETTh1, ETTh2 (Electricity Transformer Temperature datasets of the China State Grid, recorded at an hourly resolution), ETTm1, ETTm2 (15-minute variants of the same transformer monitoring system), Electricity (ECL - hourly household consumption data from 321 U.S. clients), Traffic (hourly road occupancy rates from CalTrans loop detectors in the United States), Weather (10-minute meteorological observations from 21 U.S. stations), Exchange (daily currency exchange rates across major global economies), and ILI (weekly influenza-like illness records from the U.S. CDC). These datasets are extensively used for time-series benchmarking and are publicly available [wu2022autoformer]. Forecasting horizons are set between 96, 192, 336, and 720 time steps. Long-term forecasting performance is evaluated using MSE and MAE.
The multivariate time series classification is evaluated on ten diverse datasets from the University of East Anglia (UEA) archive [bagnall2018ueamultivariatetimeseries], which includes applications from healthcare to activity recognition. To rigorously assess cross-domain transferability, we utilize seven publicly available international datasets spanning diverse applications: Human Safety/Healthcare (Fall Event [BOUBEZOUL2020105577], Risky Driving [DARUS33012023]); Human Activity Recognition (HAR) (UCI-HAR [har240], PAMAP2 [pamap2_physical_activity_monitoring_231]); Agricultural Technology/Welfare (ActBe-Calf [DISSANAYAKE2025111462]); and Industrial Monitoring/Predictive Maintenance (MetroPT-3 [metropt3], NASA Turbofan Engine [NasaTurboFan]).
2.2 Imputation
Imputation is the task of filling in missing values in a dataset. In real-world time series data, sensors can fail or data can be corrupted, which may create gaps in the data. This task tests the ability of the model to intelligently reconstruct missing data based on the surrounding context, which is vital for maintaining data quality.
As presented in Table 1, MSTN-Transformer achieves SOTA performance in 32 out of 48 ($\approx$66%) imputation evaluation scenarios and delivers the improved average results on 5 of the 6 benchmark datasets, while MSTN-BiLSTM demonstrates competitive performance. Following the TimesNet benchmark protocol [wu2023timesnettemporal2dvariationmodeling], we evaluate imputation performance on electricity and weather domain datasets, including ETT [zhou2021informer], ECL [electricityloaddiagrams20112014_321], and Weather [Kolle2025_BGCJenaWeather]. To simulate real-world scenarios, we employ random masking ratios of {12.5%, 25%, 37.5%, 50%}. This shows improved performance over several recent approaches, including GPT2(3) [zhou2023fitsallpowergeneraltime], TimesNet [wu2023timesnettemporal2dvariationmodeling], PatchTST [nie2023timeseriesworth64], LightTS [zhang2022morefastmultivariatetime], and DLinear [zeng2022transformerseffectivetimeseries].
2.3 Long-Term Forecasting
Table 2: Multivariate long-term forecasting results. Input lookback window $L=96$ ($L=36$ for ILI). Prediction horizons $H\in\{96,192,336,720\}$ ($H\in\{24,36,48,60\}$ for ILI). Avg is averaged from all four prediction lengths. Red/Blue: First/Second ranks.
Models MSTN-Trans. MSTN BiL. EMTSF LLM4TS HiMTM TIME-LLM MTST SOFTS iTransf.
Ours Ours 2025 2025 2024 2024 2024 2024 2024
H-Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
ETTm1
96 0.052 0.174 0.168 0.334 0.271 0.325 0.285 0.343 0.280 0.331 0.272 0.334 0.286 0.338 0.325 0.361 0.334 0.368
192 0.103 0.242 0.111 0.258 0.322 0.351 0.324 0.366 0.321 0.357 0.310 0.358 0.327 0.366 0.375 0.389 0.377 0.391
336 0.127 0.271 0.197 0.350 0.350 0.370 0.353 0.385 0.347 0.378 0.352 0.384 0.362 0.389 0.405 0.412 0.426 0.420
720 0.151 0.301 0.166 0.314 0.414 0.404 0.408 0.419 0.395 0.411 0.383 0.411 0.414 0.421 0.466 0.447 0.491 0.459
Avg 0.123 0.262 0.161 0.314 0.339 0.362 0.342 0.378 0.336 0.369 0.329 0.372 0.347 0.378 0.393 0.403 0.407 0.410
ETTm2
96 0.201 0.325 0.238 0.390 0.156 0.240 0.165 0.254 0.164 0.254 0.161 0.253 0.162 0.251 0.180 0.261 0.180 0.264
192 0.373 0.467 0.426 0.514 0.212 0.280 0.220 0.292 0.221 0.291 0.219 0.293 0.220 0.291 0.246 0.306 0.250 0.309
336 0.345 0.477 0.517 0.594 0.263 0.315 0.268 0.326 0.273 0.326 0.271 0.329 0.272 0.326 0.319 0.352 0.311 0.348
720 0.411 0.509 0.576 0.614 0.351 0.371 0.350 0.380 0.355 0.378 0.352 0.379 0.358 0.379 0.405 0.401 0.412 0.407
Avg 0.333 0.445 0.439 0.528 0.245 0.301 0.251 0.313 0.253 0.312 0.251 0.313 0.253 0.312 0.287 0.330 0.288 0.332
ETTh1
96 0.142 0.285 0.298 0.459 0.359 0.384 0.371 0.394 0.355 0.386 0.362 0.392 0.358 0.390 0.381 0.399 0.386 0.405
192 0.144 0.293 0.288 0.446 0.399 0.411 0.403 0.412 0.401 0.417 0.398 0.418 0.396 0.414 0.435 0.431 0.441 0.436
336 0.152 0.307 0.163 0.314 0.418 0.422 0.420 0.422 0.420 0.429 0.430 0.427 0.391 0.420 0.480 0.452 0.487 0.458
720 0.181 0.349 0.223 0.383 0.436 0.454 0.422 0.444 0.425 0.447 0.442 0.457 0.430 0.457 0.499 0.488 0.503 0.491
Avg 0.155 0.309 0.243 0.401 0.403 0.417 0.404 0.418 0.400 0.419 0.408 0.423 0.394 0.420 0.449 0.442 0.454 0.447
ETTh2
96 0.297 0.429 0.384 0.493 0.262 0.324 0.269 0.332 0.273 0.334 0.268 0.328 0.257 0.326 0.297 0.347 0.297 0.349
192 0.300 0.436 0.401 0.505 0.328 0.371 0.328 0.377 0.334 0.371 0.329 0.375 0.309 0.361 0.373 0.394 0.380 0.400
336 0.377 0.495 0.383 0.501 0.347 0.387 0.353 0.396 0.353 0.398 0.368 0.409 0.302 0.366 0.410 0.426 0.428 0.432
720 0.400 0.509 0.445 0.530 0.372 0.420 0.383 0.425 0.371 0.412 0.381 0.417 0.372 0.416 0.411 0.433 0.427 0.445
Avg 0.349 0.472 0.403 0.507 0.329 0.374 0.333 0.382 0.332 0.379 0.334 0.383 0.310 0.367 0.373 0.400 0.383 0.407
ECL
96 0.041 0.161 0.558 0.585 0.126 0.217 0.128 0.223 0.129 0.220 0.131 0.224 0.127 0.222 0.143 0.233 0.148 0.240
192 0.042 0.163 0.372 0.464 0.144 0.234 0.146 0.240 0.147 0.238 0.152 0.241 0.144 0.238 0.158 0.248 0.162 0.253
336 0.042 0.162 0.382 0.471 0.158 0.248 0.163 0.258 0.157 0.249 0.160 0.248 0.162 0.256 0.178 0.269 0.178 0.269
720 0.047 0.173 0.469 0.521 0.190 0.277 0.200 0.292 0.198 0.285 0.192 0.298 0.199 0.289 0.218 0.305 0.225 0.317
Avg 0.043 0.165 0.445 0.510 0.154 0.244 0.159 0.253 0.157 0.248 0.158 0.252 0.158 0.251 0.174 0.264 0.178 0.270
Traffic
96 0.017 0.110 0.085 0.202 0.343 0.225 0.372 0.259 0.358 0.240 0.362 0.248 0.356 0.244 0.376 0.251 0.395 0.268
192 0.018 0.111 0.086 0.205 0.369 0.238 0.391 0.265 0.368 0.248 0.374 0.247 0.375 0.251 0.398 0.261 0.417 0.276
336 0.020 0.112 0.086 0.204 0.382 0.242 0.405 0.275 0.379 0.250 0.385 0.271 0.386 0.256 0.415 0.269 0.433 0.283
720 0.020 0.113 0.087 0.207 0.424 0.270 0.437 0.292 0.430 0.276 0.430 0.288 0.425 0.279 0.447 0.287 0.467 0.302
Avg 0.019 0.111 0.086 0.205 0.379 0.243 0.401 0.273 0.384 0.254 0.388 0.264 0.386 0.257 0.409 0.267 0.428 0.282
Weather
96 0.109 0.263 0.110 0.264 0.138 0.177 0.147 0.196 0.141 0.182 0.147 0.201 0.150 0.199 0.166 0.208 0.174 0.214
192 0.111 0.264 0.111 0.265 0.181 0.220 0.191 0.238 0.188 0.228 0.189 0.234 0.194 0.240 0.217 0.253 0.221 0.254
336 0.114 0.268 0.115 0.269 0.230 0.260 0.241 0.277 0.240 0.273 0.262 0.279 0.246 0.281 0.282 0.300 0.278 0.296
720 0.128 0.286 0.132 0.289 0.304 0.315 0.313 0.329 0.312 0.322 0.304 0.316 0.319 0.333 0.356 0.351 0.358 0.347
Avg 0.115 0.270 0.117 0.272 0.213 0.243 0.223 0.260 0.220 0.251 0.225 0.257 0.227 0.263 0.255 0.278 0.258 0.278
Exchange
96 0.495 0.596 0.231 0.401 - - - - - - - - - - - - - -
192 0.456 0.566 0.278 0.437 - - - - - - - - - - - - - -
336 0.906 0.803 0.303 0.469 - - - - - - - - - - - - - -
720 0.815 0.734 0.818 0.795 - - - - - - - - - - - - - -
Avg 0.668 0.675 0.408 0.526 - - - - - - - - - - - - - -
ILI
24 0.129 0.254 0.161 0.328 1.617 0.732 - - - - 1.285 0.727 - - - - - -
36 0.155 0.287 0.167 0.333 1.586 0.728 - - - - 1.404 0.814 - - - - - -
48 0.180 0.323 0.168 0.334 1.587 0.753 - - - - 1.523 0.807 - - - - - -
60 0.190 0.332 0.169 0.334 1.560 0.768 - - - - 1.531 0.854 - - - - - -
Avg 0.165 0.302 0.166 0.332 1.588 0.745 - - - - 1.436 0.801 - - - - - -
Time series forecasting plays a crucial role in applications such as weather prediction, traffic management, and energy consumption planning. Long-term forecasting aims to predict future values far ahead in time. For example, it involves using historical weather data to estimate not only the temperature of tomorrow but also the conditions over the coming days or even weeks. We evaluate long-term forecasting performance using established protocols from the prior literature  [nie2023timeseriesworth64, wu2023timesnettemporal2dvariationmodeling, wu2022autoformer] across nine diverse benchmark datasets. The prediction horizons are set to $H\in\{96,192,336,720\}$ for the standard benchmarks and $H\in\{24,36,48,60\}$ for the ILI data set. As presented in Table 2, MSTN-Transformer and MSTN-BiLSTM show improved performance compared to specialized time series forecasting architectures, including EMTSF [alharthi2025emtsfextraordinarymixturesotamodels], LLM4TS [10.1145/3719207], HiMTM [zhao2024himtmhierarchicalmultiscalemasked], TIME-LLM [jin2024timellmtimeseriesforecasting], MTST [zhang2024multiresolutiontimeseriestransformerlongterm], SOFTS [han2024softsefficientmultivariatetime], and iTransformer [liu2024itransformerinvertedtransformerseffective].
[…]
2.4 Time-Series Classification
Table 3: Comprehensive evaluation of MSTN across diverse time-series benchmarks.
(a) Classification accuracy (%) comparison between MSTN (MSTN- Transformer and MSTN-BiLSTM) and SOTA baselines. Red/Blue: First/Second accuracy ranks.
Dataset MSTN MSTN GPT2 Times- Light DLinear Flow. ETS. FED. Stat. Auto. Py. In. Re. Trans.
Trans. BiL. (6) (Net) TS (2023) (2022) (2022) (2022a) (2022) (2021) (2021a) (2021) (2020) (2017)
(ours) (ours) (2023) (2023) (2022)
Ethanol 33.92 32.43 31.9 35.7 29.7 32.6 33.8 28.1 31.2 32.7 31.6 30.8 31.6 31.9 32.7
FaceDetect 63.20 60.30 67.3 68.6 67.5 68.0 67.6 66.3 66.0 68.0 68.4 65.7 67.0 68.6 67.3
Handwriting 43.0 35.50 32.0 32.1 26.1 27.0 33.8 32.5 28.0 31.6 36.7 29.4 32.8 27.4 32.0
Heartbeat 78.95 81.49 76.1 78.0 75.1 75.1 77.6 71.2 73.7 73.7 74.6 75.6 80.5 77.1 76.1
JapaneseV 100 100 98.6 98.4 96.2 96.2 98.9 95.9 98.4 99.2 96.2 98.4 98.9 97.8 98.7
PEMS-SF 94.32 96.59 82.1 89.6 88.4 75.1 83.8 86.0 80.9 87.3 82.7 83.2 81.5 82.7 82.1
SCP1 93.79 90.27 93.2 91.8 89.8 87.3 92.5 89.6 88.7 89.4 84.0 88.1 90.1 90.4 92.2
SCP2 62.42 60.53 59.4 57.2 51.1 50.5 56.1 55.0 54.4 57.2 50.6 53.3 53.3 56.7 53.9
SpokenArabic 99.32 99.15 99.2 99.0 100 81.4 98.8 100 100 100 100 99.6 100 97.0 98.4
UWave 76.00 68.18 88.1 85.3 80.3 82.1 86.6 85.0 85.3 87.5 85.9 83.4 85.6 85.6 85.6
(b) Generalizability Study: Performance comparison of MSTN-BiLSTM and MSTN-Transformer with TimesNet and PatchTST across seven international cross-domain datasets. Red/Blue: First/Second accuracy ranks; Cyan/Violet: First/Second inference time ranks.
MSTN TimesNet PatchTST
Dataset Det. Dom. Cnt. A.Block Acc.(%) F1 Prec. Rec. Size I.Time Acc.(%) I.Time Acc.(%) I.Time Prior Acc.
Rodegast[DARUS33012023] Sim. Hum. DE. BiLSTM 99.20 0.992 0.993 0.991 4.39 0.81 99.18 5.07 98.64 4.80 91.00[RF,GB][Rodegast2024]
Transf. 99.53 0.947 0.951 0.942 0.54 0.155
Boubezoul[BOUBEZOUL2020105577] Real. Hum. FR. BiLSTM 93.75 0.945 0.940 0.950 4.16 2.79 93.00 4.20 91.37 4.68 91.59[DT][fi15100333]
Transf. 91.76 0.927 0.921 0.934 0.96 0.24
UCI-HAR[har240] Act. Hum. IT. BiLSTM 96.59 0.965 0.966 0.965 7.14 2.03 91.38 45.30 93.21 4.36 83.35[Km,NB][Ismi2016KmeansCB]
Transf. 95.58 0.956 0.958 0.955 29.98 3.48
PAMAP2[pamap2_physical_activity_monitoring_231] Phy. Hum. US. BiLSTM 99.69 0.996 0.996 0.996 0.76 0.086 95.13 0.46 98.02 3.21 90.00[kNN][6246152]
Transf. 99.52 0.995 0.995 0.995 1.58 0.20
ActBeC.[DISSANAYAKE2025111462] Calf. Anim. IE. BiLSTM 93.00 0.923 0.923 0.923 4.14 1.76 90.65 8.53 62.45 2.20 84.00[RCCV][DISSANAYAKE2025111462]
Transf. 88.14 0.883 0.873 0.913 0.89 0.14
MetroPT3[metropt3] Met. Mech. PT. BiLSTM 93.10 0.930 0.930 0.931 0.57 0.042 93.00 0.09 81.67 0.19 62.00[SAE][9564181]
Transf. 93.83 0.938 0.940 0.938 1.30 0.19
NASA[NasaTurboFan] Eng. Mech. US. BiLSTM 20.24 4.25 1.48 15.56 4.52 31.59 2.31
Transf. 11.26 1.31 0.45
Abbreviations: Det: Details (Sim: Simulator PTW data, Real: Real PTW Data, Act: Activity human, Phy: Physical activity, Calf: Calf Behavior, Met: Metro predictive maintenance, Eng: Engine sensor), Dom: Domain (Hum: Human Safety, Anim: Animal welfare, Mech: Mechanical), Cnt: Country (IN: India, FR: France, DE: Germany, IT: Italy, US: USA, IE: Ireland, PT: Portugal), A.Block: Architecture block (Transf. Transformer), I.Time: Inference time, †RMSE values.
Classification involves categorizing entire time series sequences. It is used to identify which type of event or pattern the sequence represents. For example, this could be the classification of a segment of sensor data as walking, running, falling, or collision activity, which is fundamental for automated monitoring and diagnosis systems. Performance is evaluated using the standard metric of classification accuracy.
Table 3 presents the classification results across 10 UEA standard benchmark datasets. The proposed MSTN framework is compared against SOTA time-series classification methods, including TimesNet [wu2023timesnettemporal2dvariationmodeling], GPT-2(6) [zhou2023fitsallpowergeneraltime], FEDformer [zhou2022fedformer], and LightTS [zhang2022morefastmultivariatetime]. MSTN demonstrates competitive performance in multiple datasets. MSTN-Transformer achieves perfect classification (100%) on JapaneseVowels and near-perfect performance on SpokenArabic (99.32%), while MSTN-BiLSTM achieves improved accuracy on Heartbeat (81.49%), JapaneseVowels (100%) and PEMS-SF (96.59%).
2.5 Generalizability Study
A robust model should not be a specialist in just one area; it should adapt to new tasks and data distributions across different data and domains. This experiment evaluates the ability of the MSTN to transfer its knowledge across diverse domains, from human activity recognition to mechanical fault prediction. As summarized in Table 3, the proposed MSTN framework demonstrates generalization in seven international benchmark data sets such as human safety, activity recognition, animal welfare, and mechanical prognostics. Without domain-specific tuning, MSTN shows improved performance over the latest time series SOTA models, including TimesNet [wu2023timesnettemporal2dvariationmodeling] and PatchTST [nie2023timeseriesworth64].
MSTN achieves competitive performance across fundamentally different domains, demonstrating its capabilities as a universal temporal model. In human safety applications, MSTN-Transformer reaches 99.53% accuracy on German collision prediction data [DARUS33012023], substantially outperforming prior work (91.0% RF,GB [Rodegast2024]). For activity recognition, MSTN-BiLSTM achieves 96.59% accuracy on the UCI-HAR dataset [har240], representing a 13.2% improvement over TimesNet. The framework also excels in real-world fall detection with 93.75% accuracy, surpassing dedicated DT-based methods. MSTN-BiLSTM also shows strong performance in animal welfare evaluation [DISSANAYAKE2025111462], assessment with 93.00% accuracy on calf behavior recognition. Furthermore, MSTN-Transformer demonstrates better capability in mechanical prognostics, achieving 11.26 RMSE on the NASA Turbofan data set [NasaTurboFan] and significantly outperforming all baseline predictors in the estimate of the remaining useful life.
[…]
2.6 Ablation Study
Table 4: Ablation study of MSTN-Transformer and MSTN-BiLSTM. $H\in\{96,192,336,720\}$ and for ILI $H\in\{24,36,48,60\}$. Red and blue indicate first and second best results, respectively.
Methods MSTN Full w/o CNN w/o Transf. w/o SE w/o MHA w/o SGF
Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
MSTN Transformer
ECL
96 0.041 0.161 0.045 0.165 0.048 0.173 0.042 0.163 0.043 0.164 0.045 0.168
192 0.042 0.163 0.046 0.169 0.048 0.173 0.045 0.166 0.044 0.166 0.047 0.170
336 0.042 0.162 0.046 0.168 0.050 0.178 0.046 0.167 0.045 0.169 0.049 0.168
720 0.047 0.173 0.051 0.179 0.055 0.186 0.049 0.175 0.050 0.178 0.053 0.179
Weather
96 0.109 0.263 0.112 0.266 0.118 0.273 0.110 0.264 0.111 0.265 0.112 0.265
192 0.111 0.264 0.115 0.269 0.120 0.275 0.112 0.265 0.113 0.265 0.112 0.265
336 0.114 0.268 0.120 0.274 0.123 0.279 0.115 0.269 0.116 0.270 0.116 0.270
720 0.128 0.286 0.138 0.296 0.136 0.294 0.130 0.287 0.131 0.289 0.130 0.287
ETTm1
96 0.052 0.174 0.062 0.186 0.089 0.227 0.061 0.185 0.057 0.178 0.055 0.177
192 0.104 0.242 0.106 0.244 0.105 0.243 0.109 0.247 0.103 0.244 0.105 0.245
336 0.127 0.271 0.146 0.289 0.156 0.307 0.140 0.283 0.151 0.296 0.147 0.290
720 0.151 0.301 0.155 0.306 0.172 0.319 0.142 0.288 0.159 0.308 0.158 0.307
ETTh2
96 0.297 0.429 0.370 0.492 0.385 0.498 0.374 0.482 0.377 0.491 0.390 0.485
192 0.300 0.436 0.404 0.511 0.391 0.505 0.392 0.506 0.382 0.496 0.395 0.501
336 0.377 0.495 0.415 0.520 0.410 0.518 0.396 0.508 0.387 0.499 0.396 0.505
720 0.400 0.509 0.428 0.524 0.430 0.525 0.426 0.523 0.426 0.531 0.439 0.531
Traffic
96 0.017 0.110 0.020 0.113 0.026 0.128 0.019 0.112 0.021 0.116 0.020 0.112
192 0.018 0.111 0.020 0.115 0.026 0.127 0.020 0.113 0.023 0.116 0.020 0.113
336 0.020 0.112 0.022 0.118 0.029 0.135 0.021 0.115 0.024 0.118 0.021 0.114
720 0.020 0.113 0.026 0.128 0.026 0.128 0.022 0.118 0.025 0.119 0.025 0.120
ILI
24 0.129 0.254 0.140 0.272 0.139 0.274 0.131 0.255 0.131 0.255 0.133 0.256
36 0.155 0.287 0.167 0.301 0.168 0.307 0.161 0.298 0.157 0.293 0.168 0.304
48 0.180 0.323 0.208 0.344 0.193 0.332 0.182 0.324 0.183 0.325 0.191 0.328
60 0.190 0.332 0.197 0.338 0.199 0.339 0.193 0.333 0.194 0.334 0.194 0.338
MSTN BiLSTM
Methods MSTN Full w/o CNN w/o BiLSTM w/o SE w/o MHA w/o SGF
Weather
96 0.110 0.264 0.112 0.266 0.119 0.274 0.111 0.264 0.111 0.265 0.112 0.266
192 0.111 0.265 0.113 0.267 0.120 0.275 0.112 0.266 0.113 0.267 0.113 0.266
336 0.115 0.269 0.116 0.270 0.123 0.279 0.116 0.270 0.117 0.271 0.118 0.272
720 0.132 0.289 0.138 0.295 0.139 0.296 0.134 0.294 0.135 0.294 0.134 0.298
ETTm2
96 0.238 0.390 0.243 0.400 0.264 0.430 0.262 0.394 0.240 0.392 0.261 0.399
192 0.426 0.514 0.501 0.560 0.528 0.617 0.408 0.510 0.419 0.535 0.411 0.512
336 0.517 0.594 0.581 0.623 0.571 0.612 0.535 0.592 0.536 0.593 0.626 0.655
720 0.576 0.614 0.660 0.651 0.671 0.665 0.638 0.645 0.829 0.738 0.641 0.647
Traffic
96 0.085 0.202 0.089 0.208 0.090 0.209 0.086 0.203 0.087 0.204 0.087 0.205
192 0.087 0.205 0.092 0.218 0.101 0.221 0.088 0.207 0.088 0.208 0.088 0.209
336 0.086 0.204 0.095 0.224 0.105 0.227 0.089 0.213 0.090 0.214 0.090 0.215
720 0.087 0.207 0.098 0.227 0.107 0.229 0.091 0.218 0.091 0.217 0.093 0.218
An ablation study systematically removes individual components from the full model to evaluate their contributions. Table 4 presents the results for MSTN-Transformer across six diverse benchmarks ECL (electricity), Weather, ETTm1, ETTh2, Traffic, ILI and for MSTN-BiLSTM on three datasets: Weather, ETTm2, and Traffic. For MSTN-Transformer, five variants are tested: w/o CNN, w/o Transformer, w/o SE, w/o MHA, and w/o SGF. While In MSTN-BiLSTM architecture, w/o BiLSTM denotes the variant of MSTN-BiLSTM without the BiLSTM layer.
In the MSTN-Transformer variant, the results consistently show that the full MSTN-Transformer model achieves the best performance (ranked first in MSE/MAE) across nearly all prediction horizons and datasets. Removing the core Transformer module (w/o Transformer) leads to the most severe performance drop in most cases. For instance, on ETTh2 (a challenging dataset with high volatility), the MSE increases substantially from 0.297 (full model) to 0.385 at horizon 96 when the Transformer is ablated, highlighting its critical role in modeling complex temporal patterns. Similarly, the CNN pathway (w/o CNN) is highly important, especially on datasets like Traffic and ECL where local features are crucial; its removal causes noticeable degradation (e.g., Traffic H=96 MSE rises from 0.017 to 0.020). The SE blocks (w/o SE) also contribute significantly, often yielding the second-best or competitive results when present. For example, on Weather, the SE-ablated variant remains close to the full model, but on ETTm1, its removal increases MSE from 0.127 to 0.140 at H=336. The MHA (w/o MHA) and self-gated fusion (w/o SGF) modules further enhance robustness, with their absence leading to consistent, though sometimes smaller, declines across datasets.
For MSTN-BiLSTM, the full model outperforms all ablated variants. Removing the BiLSTM layer (w/o BiLSTM) results in significant deterioration, particularly on complex datasets like ETTm2 where MSE at H=720 jumps from 0.576 to 0.671. This underscores the importance of recurrent modeling in capturing sequential dependencies. In summary, the ablation study validates that each architectural component is essential. The Transformer or BiLSTM, convolutional pathway, and attention-based mechanisms operate synergistically, with the full integrated model delivering the best overall forecasting accuracy across diverse temporal datasets.
2.7 Model Complexity and Edge-AI Deployability
2.7.1 Complexity Analysis
In this analysis, we denote the length of the lookback window as $L$, the number of variate features (channels) as $C$, and the task-specific output dimension as $K$. We use $C$ for channels to align with notation in common literature such as SOFTS [han2024softsefficientmultivariatetime], where $K$ is often referred to as the forecasting horizon ($H$). We simplify the asymptotic complexity expressions by omitting the model dimension $d$, assuming $d\approx C$.
The MSTN framework addresses the fundamental trade-off in temporal modeling between representational capacity and computational efficiency. Although the proposed variants exhibit distinct encoding complexities with MSTN-BiLSTM maintaining strict linearity with respect to sequence length and MSTN-Transformer incurring a quadratic cost. They both converge on a unified efficiency strategy: ETA. This is achieved because both the parallel CNN and sequence modeling pathways utilize immediate temporal pooling to eliminate sequence-length dependence in subsequent layers. Specifically, both the parallel CNN and sequence modeling pathways (Transformer/BiLSTM) independently condense their temporal features via Global Average Pooling and Sequence Mean Pooling, respectively (reducing the sequence dimension from $L$ to $1$). Following concatenation, the subsequent SGF, SE-Block, and MHA refinement stages operate on static feature vectors. This strategic design ensures that the refinement pipeline scales as $\mathcal{O}(C^{2})$ with respect to channels but remains constant-time ($\mathcal{O}(1)$) with respect to the input length $L$, effectively isolating the heavy temporal computation to the initial stage.
Component-Wise Analysis of MSTN-Transformer: The complexity is dominated by the sequence encoding stage. First, the Multi-Scale CNN extracts local features with linear complexity $\mathcal{O}(CL)$. Second, the Transformer Encoder applies self-attention; as this requires computing an $L\times L$ attention matrix, it introduces the dominant quadratic term $\mathcal{O}(CL^{2})$. Following this, Temporal Pooling aggregates the sequence with linear cost $\mathcal{O}(CL)$. Crucially, the Refinement stage operates on the pooled vector, scaling with channel interactions as $\mathcal{O}(C^{2})$. Finally, the Projection Head maps features to the output with complexity $\mathcal{O}(CK)$. Summing these components, the total asymptotic complexity is $\mathcal{O}(CL^{2}+C^{2}+CK)$, which scales quadratically with sequence length $L$.
Component-Wise Analysis of MSTN-BiLSTM: In contrast, the MSTN-BiLSTM variant achieves strict linear efficiency with respect to sequence length. The BiLSTM Encoder processes the sequence recurrently without generating a global attention map, resulting in a linear complexity of $\mathcal{O}(CL)$. Similar to the Transformer variant, Temporal Pooling ($\mathcal{O}(CL)$) condenses the time dimension. The Refinement stage contributes $\mathcal{O}(C^{2})$, and the Projection Head contributes $\mathcal{O}(CK)$. Since the encoder avoids quadratic operations in $L$, the total start-to-end complexity remains highly efficient at $\mathcal{O}(CL+C^{2}+CK)$, which scales linearly with $L$.
Table 5: Comprehensive evaluation of MSTN across diverse time-series benchmarks. Inference time of MSTN compared to models including SOFTS, iTransformer, TimesNet and PatchTST. Red/Blue: First/Second ranks.
(a) Comparison of MSTN with the MSE and inference time (I.T) metrics reported in SOFTS [han2024softsefficientmultivariatetime] on the Traffic dataset ($L=96$, $H=720$).
MSE Perf. Gap I.T.(ms) Speed Gap
[…]
(b) Comprehensive evaluation of MSTN across diverse time-series forecasting benchmarks [wu2022autoformer].
Dataset Horizon Size (MB) I.T. (ms)
Traffic 96,192,336,720 2.87 0.72
ETTh2 96,192,336,720 3.65 0.91
ETTm2 96,192,336,720 2.92 0.85
Weather 96,192,336,720 2.85 0.70
Exchange 96,192,336,720 2.96 0.80
ILI 24,36,48,60 2.51 0.63
(c) Inference time (I.T., in ms) of MSTN variants compared to TimesNet and PatchTST across seven international datasets.
Dataset MSTN Size (MB) I.T. TimesNet I.T. PatchTST I.T.
Rodegast [DARUS33012023] MSTN-Trans. 0.54 0.15 5.07 4.80
MSTN-BiLSTM 4.39 0.81
Boubezoul [BOUBEZOUL2020105577] MSTN-Trans. 0.96 0.24 4.20 4.68
MSTN-BiLSTM 4.16 2.79
UCI-HAR [har240] MSTN-Trans. 29.98 3.48 45.30 4.36
MSTN-BiLSTM 7.14 2.03
PAMAP2 [pamap2_physical_activity_monitoring_231] MSTN-Trans. 1.58 0.20 0.46 3.21
MSTN-BiLSTM 0.76 0.086
ActBeCalf [DISSANAYAKE2025111462] MSTN-Trans. 0.89 0.14 8.53 2.20
MSTN-BiLSTM 4.14 1.76
MetroPT3 [metropt3] MSTN-Trans. 1.30 0.19 0.09 0.19
MSTN-BiLSTM 0.57 0.042
NASA [NasaTurboFan] MSTN-Trans. 1.31 0.45 4.52 2.31
MSTN-BiLSTM 4.25 1.48
2.7.2 Comparative Analysis and Edge-AI Deployability
Recent efficient architectures demonstrate varying complexity profiles: linear models such as DLinear [zeng2022transformerseffectivetimeseries] achieve $\mathcal{O}(CL)$ complexity, while spectral approaches like FEDformer [zhou2022fedformer] and TimesNet [wu2023timesnettemporal2dvariationmodeling] achieve $\mathcal{O}(CL\log L)$ complexity. Following the notation in SOFTS [han2024softsefficientmultivariatetime], their model achieves $\mathcal{O}(CL+CH)$ complexity via stochastic pooling, while iTransformer ($\mathcal{O}(C^{2}+CL+CH)$), PatchTST ($\mathcal{O}(CL^{2}+CH)$), and Standard Transformers ($\mathcal{O}(N\cdot CL^{2})$) incur heavier costs.
Against this landscape, the proposed MSTN-BiLSTM achieves strict linear temporal complexity $\mathcal{O}(CL+C^{2}+CK)$. Theoretically, this matches the optimal efficiency of SOFTS in the predominant real-world scenario where sequence length exceeds channel dimension ($L\gg C$). While SOFTS retains a theoretical advantage in high-dimensional, short-sequence settings ($C\gg L$) due to its purely linear channel dependence, such data structures are rare in long-term time series forecasting tasks. Meanwhile, MSTN-Transformer scaling as $\mathcal{O}(CL^{2}+C^{2}+CK)$ optimizes practical runtime by confining the quadratic term to a single encoder layer rather than a deep stack ($N$ layers) as seen in iTransformer or standard Transformer architectures. This strategic design allows MSTN to balance the high-capacity modeling of Transformers with the latency requirements of Edge-AI, delivering state-of-the-art performance with minimized computational overhead.
One thing to note is that despite the MSTN-BiLSTM’s better theoretical scaling ($\mathcal{O}(L)$) than the MSTN-Transformer’s $\mathcal{O}(L^{2})$, the latter runs significantly faster than the former (Table 5). This paradox is due to the fundamental conflict between the model’s structure and the way modern acceleration architectures process the data. The BiLSTM encoder is inherently recurrent, meaning that the calculation of the hidden state $h_{t}$ is strictly dependent on the previous state $h_{t-1}$. This sequential data dependency prevents the hardware’s parallel processing units from being fully utilized, forcing the computation into a slow, iterative loop (low hardware utilization). In contrast, the Transformer executes its $\mathcal{O}(L^{2})$ self-attention via highly optimized [vaswani2023attentionneed], complete matrix parallelization. The core operation, the $QK^{T}$ matrix multiplication, is performed simultaneously across thousands of processing elements. This high utilization of the hardware’s parallel capacity, confined to a single encoder layer before pooling, ensures the Transformer’s execution speed drastically outweighs the theoretical linearity of the BiLSTM’s sequential cost, resulting in lower practical latency across high-throughput platforms.
Standard Transformers maintain the full sequence length $L$ across a multi-layer stack of $N$ blocks for self-attention operations, resulting in cumulative attention cost proportional to $N\cdot\mathcal{O}(CL^{2})$.  In contrast, MSTN applies attention in its encoder, then immediately pools ($L\to 1$). Standard Transformers maintain $L$ across all layers, while MSTN eliminates $L$-dependence after encoding. Consequently, the remainder of the network operates on static feature vectors with $\mathcal{O}(1)$ complexity relative to $L$ for operations involving sequence length. By eliminating the recursive application of attention across deep layers, MSTN reduces the sequence-dependent FLOPs significantly. While MSTN-BiLSTM avoids quadratic attention costs entirely by employing BiLSTM encoding with $\mathcal{O}(CL)$ complexity. When combined with the remaining pipeline components (CNN, refinement, and projection), the total complexity is $\mathcal{O}(CL+C^{2}+CK)$, which scales linearly with sequence length $L$.
Following the baseline results and inference time metrics reported in SOFTS [han2024softsefficientmultivariatetime] on the Traffic data set (lookback window $L=96$, horizon $H=720$), we compare our MSTN results with their metrics reported in Table 5. MSTN-Transformer demonstrates that architectural efficiency can surpass asymptotic predictions. MSTN-Transformer achieves $69\times$ faster inference than PatchTST (0.72 ms vs. 50.00 ms). Furthermore, it outperforms models with better theoretical scaling, showing speedups of $34.7\times$ over SOFTS and $305\times$ over TimesNet. This practical gain, irrespective of the asymptotic class, is primarily delivered by the early temporal pooling of the MSTN architecture ($L\to 1$), a strategic design choice that dramatically reduces the overall FLOPs compared to deep, unpooled attention stacks. This performance represents a $22.4\times$ accuracy improvement and $34.7\times$ speedup over the SOFTS baseline, and a significant speed and accuracy advantage over all other models detailed in Table 5. This improved efficiency enables substantially reduced memory usage and inference latency, positioning MSTN as a practical solution for deployment on constrained hardware platforms.
As summarized in Table 5, MSTN demonstrates consistently efficient inference across six widely used long-term forecasting benchmarks. Despite the variation in dataset scale and temporal structure, MSTN maintains a compact model size in the range of 2.5–3.7 MB and sustains sub-millisecond latency across all horizons, including the longer 336 and 720-step settings. This stability in runtime behavior across heterogeneous datasets such as Traffic, ETTh2, ETTm2, Weather, Exchange, and ILI highlights the robustness of MSTN’s dual-path, multi-scale feature extraction. The uniformly low inference cost emphasizes its suitability for real-time forecasting scenarios and deployment in resource-constrained environments where computational consistency is critical.
[…]
3 Discussion
This work introduces MSTN, a DL framework founded on a hierarchical multi-scale modeling principle and defined by its strategic use of ETA. MSTN integrates core components: a multi-scale convolutional encoder that constructs a hierarchical feature pyramid; a sequential modeling component for long-term global dependency modeling, empirically validated with BiLSTM and Transformer variants; and a SGF mechanism augmented with SE and MHA to enable dynamic, context-aware feature integration. Existing methods often rely on fixed-scale structural priors (e.g., patch-based tokenization, fixed frequency transformations, or frozen backbone architectures). These limitations have historically forced a compromise between scale diversity and inference speed. In contrast, our approach efficiently integrates convolutional local pattern extraction with long-term global dependency modeling, leveraging ETA via an adaptive SGF mechanism. This enables our model to capture multi-scale temporal features while maintaining computational efficiency and simultaneously achieving SOTA performance.
[…]
Although MSTN excels in standard forecasting regimes ($L\gg C$), we acknowledge that factorization-based models such as SOFTS retain a theoretical efficiency advantage in high-dimensional, short-sequence settings ($C\gg L$). However, since such datasets are rarely encountered in long-term time series forecasting tasks, this limitation has a negligible impact on practical deployability.
Future work will explore cross-variate attention mechanisms and large-scale pre-training to further enhance the capabilities of MSTN as a foundation model for time series understanding. Our work paves the way for efficient multi-scale temporal modeling under resource constraints, enabling real-time applications across diverse domains from healthcare to industrial monitoring.
4 Methods
This section details the MSTN architecture and its methodological innovations, followed by the baseline models, and training configurations.
4.1 Proposed Architecture: MSTN
4.1.1 Architecture Overview
(a) MSTN architecture
(b) MSTN multi-scale signal processing pipeline
Fig. 1: Proposed MSTN: (a) architectural diagram and (b) signal processing pipeline
The proposed Multi-Scale Temporal Network (MSTN; Fig. 1a) is a hybrid deep-learning architecture designed around a parallel multi-scale modeling paradigm. MSTN processes temporal signals through two coordinated encoding pathways, each specialized for complementary aspects of temporal structure. A key design element is the use of an early temporal aggregation (ETA) mechanism, applied directly to the outputs of the dual encoders, which collapses the full input sequence of length $L$ into a single fixed-dimensional representation ($L\rightarrow 1$). This operation ensures that all subsequent refinement stages operate in constant time with respect to sequence length, thereby enabling efficient processing of long temporal inputs.
The network comprises three principal modules. (i) A multi-scale convolutional branch captures localized temporal patterns using hierarchically stacked convolutions; its output is aggregated via global average pooling, promoting robustness and translation invariance. (ii) A sequence modeling branch captures long-range dependencies, instantiated using either a bidirectional LSTM (MSTN-BiLSTM) or a Transformer encoder (MSTN-Transformer). To maintain computational efficiency, this branch is compressed through sequence mean pooling, yielding a global summary vector. (iii) A multi-scale fusion and refinement module integrates the pooled representations from the two branches. The features are concatenated and passed through a scale-gated fusion (SGF) mechanism, which adaptively modulates the contributions of each pathway. The fused representation is subsequently refined using channel-wise recalibration via squeeze–excitation blocks and multi-head attention, enabling the network to emphasize the most informative feature dimensions. The final representation is normalized, regularized through dropout, and projected to the task-specific output space. To address class imbalance, training is performed using a focal loss objective. Figure 1b provides an overview of the end-to-end processing pipeline, from parallel feature extraction to multi-scale fusion and refinement.
4.1.2 Methodology
We consider the following problem: Given multivariate time series samples $X_{1:L}\in\mathbb{R}^{L\times C}$ with a lookback window sequence length $L$ and $C$ variate features, MSTN learns temporal representations through a novel parallel architecture. We detail the complete process below.
Forward Process. The input $X\in\mathbb{R}^{B\times L\times C}$ is processed through dual complementary pathways that operate simultaneously. The CNN pathway captures local temporal patterns, while the sequence modeling pathway (implemented as either a Transformer or a BiLSTM) captures long-range temporal dependencies. Their outputs are fused through a sophisticated gating and attention mechanism for enhanced representation learning. The input undergoes successive 1D convolutions to extract multi-scale features, adjusting the feature dimension while preserving temporal resolution, followed by refinement of the representations. The $\text{Conv1D}_{k}$ convolutional layers employ a hierarchical design where the first layer with kernel size $k=7$ extracts basic temporal patterns, and the second layer with kernel size $k=5$ learns combinations of these patterns into more complex features, with padding to maintain temporal dimensions, and batch normalization stabilizes training. To aggregate these temporal features into a compact representation, we apply global average pooling across the entire sequence:
$H_{\text{conv}}^{(1)}$ $=\text{ReLU}(\text{BN}(\text{Conv1D}_{7}(X^{T})))\in\mathbb{R}^{B\times 128\times L}$ (1)
$H_{\text{conv}}^{(2)}$ $=\text{ReLU}(\text{Conv1D}_{5}(H_{\text{conv}}^{(1)}))\in\mathbb{R}^{B\times 64\times L}$ (2)
$\mathbf{z}_{\text{cnn}}$ $=\text{GlobalAvgPool1D}(H_{\text{conv}}^{(2)})\in\mathbb{R}^{B\times 64}$ (3)
where each of the 64 feature channels is summarized by its average activation over time through temporal pooling.
We employ two sequence modeling approaches: MSTN-Transformer and MSTN-BiLSTM. In MSTN-Transformer, the architecture captures long-range dependencies through a single Transformer encoder block composed of 4 layers with 8 attention heads. The input is processed through self-attention mechanisms where each of the L time steps maintains a 128-dimensional representation enriched with global contextual information. To consolidate these temporal representations, we apply sequence mean pooling across the time dimension:
$H_{\text{trans}}=\text{TransformerEnc}(X)\in\mathbb{R}^{B\times L\times 128};\quad\mathbf{z}_{\text{trans}}=\frac{1}{L}\sum_{t=1}^{L}H_{\text{trans},t}\in\mathbb{R}^{B\times 128}$ (4)
In MSTN-BiLSTM: The architecture captures bidirectional sequential dependencies, processing sequences in both forward and backward directions to learn complex temporal dynamics and memorize long-range patterns:
$H_{\text{bilstm}}=\text{BiLSTM}(X)\in\mathbb{R}^{B\times L\times 128};\quad\mathbf{z}_{\text{bilstm}}=\frac{1}{L}\sum_{t=1}^{L}H_{\text{bilstm},t}\in\mathbb{R}^{B\times 128}$ (5)
The BiLSTM employs 2 layers with 64 hidden units per direction, capturing both causal and anti-causal temporal relationships.
The CNN pathway utilizes global average pooling, and the sequence modeling pathways employ sequence mean pooling. These pooling operations, along with the subsequent reshaping, collectively form the ETA module, which collapses the temporal dimension $L$ to $1$. This is the critical step that ensures subsequent operations happen at an $\mathcal{O}(1)$ inference cost. The parallel features undergo sophisticated fusion through a multi-stage enhancement process that combines the strengths of both pathways. For MSTN-Transformer:
$\mathbf{z}_{\text{concat}}=[\mathbf{z}_{\text{cnn}};\mathbf{z}_{\text{trans/bilstm}}]\in\mathbb{R}^{192}$ (6)
where $\mathbf{z}_{\text{cnn}}\in\mathbb{R}^{64}$ and $\mathbf{z}_{\text{trans}}\in\mathbb{R}^{128}$. The architecture then performs adaptive weighting through an SGF mechanism, illustrated in Fig. 1, that learns to dynamically balance feature contributions:
$\mathbf{z}_{\text{fused}}=\mathbf{z}_{\text{concat}}\odot\sigma(W_{g}\mathbf{z}_{\text{concat}}+b_{g})$ (7)
where for both variants: $W_{g}\in\mathbb{R}^{192\times 192}$, $b_{g}\in\mathbb{R}^{192}$. This mechanism learns to emphasize the most relevant features from each pathway, creating a harmonious blend of multi-scale temporal features. The fused features are reshaped for efficient single-token processing: $\mathbf{z}_{\text{seq}}=\mathbf{z}_{\text{fused}}\otimes\mathbf{1}\in\mathbb{R}^{B\times 1\times 192}$. Channel-wise attention then enhances features through an SE block. Given the temporal dimension is already $L=1$, the global pooling step within SE is trivially executed: $\mathbf{z}_{\text{se}}=\mathbf{z}_{\text{seq}}\odot\sigma(W_{2}\text{ReLU}(W_{1}\mathbf{z}_{\text{seq}}))$ where $W_{1}\in\mathbb{R}^{24\times 192}$, $W_{2}\in\mathbb{R}^{192\times 24}$ with reduction ratio 8. The SE mechanism amplifies informative channels while suppressing less useful ones. Feature dependencies are refined through MHA. Since the temporal dimension is condensed ($L=1$), the MHA block functions here as a sophisticated global feature recalibration layer rather than a temporal mixer, confirming $\mathcal{O}(1)$ complexity:
$\mathbf{z}_{\text{final}}=\text{Dropout}\left(\text{LayerNorm}\left(\sum_{t=1}^{1}\text{MHA}(\mathbf{z}_{\text{se}})_{t}\right),p=0.3\right)\in\mathbb{R}^{B\times 192}$ (8)
where 192 is the feature dimension for both MSTN-Transformer and MSTN-BiLSTM, with 4 attention heads (dim=48 per head). The refined features $\mathbf{z}_{\text{final}}$ are then passed to the final task-specific head (imputation, forecasting, or classification). The refined features $\mathbf{z}_{\text{final}}$ are then passed to the final task-specific head (imputation, forecasting, or classification). The prediction head consists of a single linear layer that maps the $192$-dimensional aggregated feature vector to the task-specific output shape.
Task-Specific Prediction and Loss Functions: We detail the prediction head ($\hat{Y}$ or $\hat{X}$) and the corresponding loss function $\mathcal{L}$ for each of the three tasks:
Time Series Imputation: The head reconstructs the input sequence $\hat{X}_{1:L}\in\mathbb{R}^{B\times L\times C}$, and the loss is calculated as masked MSE only over the observed values ($\Omega$ is the mask matrix):
$\hat{X}_{1:L}$ $=W_{i}\mathbf{z}_{\text{final}}+b_{i}\in\mathbb{R}^{B\times L\times C}$ (9)
$\mathcal{L}_{\text{imputation}}$ $=\frac{1}{|\Omega|}\sum_{i=1}^{B}\sum_{t=1}^{L}\sum_{c=1}^{C}\Omega_{i,t,c}\cdot\left(\hat{X}_{i,t,c}-X_{i,t,c}\right)^{2}$ (10)
Long-Term Forecasting: The head produces a forecast $\hat{Y}_{1:T}\in\mathbb{R}^{B\times T\times C}$ for a horizon $T$, and the objective is to minimize the MSE:
$\hat{Y}_{1:T}$ $=W_{f}\mathbf{z}_{\text{final}}+b_{f}\in\mathbb{R}^{B\times T\times C}$ (11)
$\mathcal{L}_{\text{forecast}}$ $=\frac{1}{2TC}\sum_{i=1}^{B}\sum_{t=1}^{T}\sum_{c=1}^{C}\left(\hat{Y}_{i,t,c}-Y_{i,t,c}\right)^{2}$ (12)
Time Series Classification: The head computes class probabilities $P\in\mathbb{R}^{B\times K}$ over $K$ classes, and the loss is the canonical Cross-Entropy (CE) Loss:
$P$ $=\text{Softmax}(W_{c}\mathbf{z}_{\text{final}}+b_{c})\in\mathbb{R}^{B\times K}$ (13)
$\mathcal{L}_{\text{classify}}$ $=-\frac{1}{B}\sum_{i=1}^{B}\sum_{k=1}^{K}Y_{i,k}\log(P_{i,k})$ (14)
where $Y$ is the one-hot encoded true label. This structure allows the MSTN framework to be universally applied to various multivariate time series tasks.
4.2 Method Level Innovation
Table 6: Analysis of model architectures by design characteristics. This table outlines how different state-of-the-art models align with the established Channel Interaction (CI, CD, CP) strategies based on the temporal and channel characteristics they model. MSTN is positioned as a hybrid design (CI+CD). (Asym.: Asymmetry, Lag.: Lagginess, Pol.: Polarity, Gw.: group-wise, Dyn.: Dynamism, Ms.: Multi-scale)
Strategy Mechanism Characteristic Method
Asym. Lag. Pol. Gw. Dyn. Ms.
CI - - - - - - - PatchTST
CI - - - - - - - CycleNet
CI - - - - - - - DLinear
CI - - - - - - - Timer
CI - - - - - - - Chronos
CI - - - - - - - LLM4TS
CI - - - - - - - Time-LLM
CI - - - - - - - RevIN
CD CNN-based $\checkmark$ - - - - - Informer
CD CNN-based $\checkmark$ - - - - - Autoformer
CD CNN-based $\checkmark$ - - - - - FEDformer
CD CNN-based $\checkmark$ - - - - - TimesNet
CD MLP-based $\checkmark$ - - - - - TSMixer
CD MLP-based $\checkmark$ - - - - - TTM
CD Transformer-based $\checkmark$ - - - $\checkmark$ - iTransformer
CD Transformer-based $\checkmark$ - - - $\checkmark$ - Crossformer
CD Transformer-based $\checkmark$ $\checkmark$ - - - - VCformer
CD Transformer-based $\checkmark$ $\checkmark$ - - $\checkmark$ - MOIRAI
CD Transformer-based $\checkmark$ - - - $\checkmark$ - UniTS
CD GNN-based - - - $\checkmark$ - - GTS
CD GNN-based $\checkmark$ - - - - $\checkmark$ MSGNet
CD GNN-based - $\checkmark$ - - - - FourierGNN
CD GNN-based - $\checkmark$ - - - - FC-STGNN
CD GNN-based $\checkmark$ - - - $\checkmark$ - TPGNN
CD Linear Channel Mixing (Factorization) $\checkmark$ - - - - - SOFTS
CD Others $\checkmark$ - - - $\checkmark$ - C-LoRA
CP CNN-based $\checkmark$ - - - - - ModernTCN
CP Transformer-based $\checkmark$ - - $\checkmark$ - - DUET
CP Transformer-based $\checkmark$ - - - $\checkmark$ - MCformer
CP Transformer-based $\checkmark$ - - $\checkmark$ $\checkmark$ - DGCformer
CP Transformer-based - - - - $\checkmark$ - CM
CP GNN-based $\checkmark$ - - - - - MTGNN
CP GNN-based $\checkmark$ - $\checkmark$ - - - CrossGNN
CP GNN-based $\checkmark$ - - $\checkmark$ - - WaveForM
CP GNN-based - - - - $\checkmark$ - MTSF-DG
CP GNN-based $\checkmark$ - - $\checkmark$ - - ReMo
CP GNN-based $\checkmark$ - - $\checkmark$ $\checkmark$ $\checkmark$ Ada-MSHyper
CP Others $\checkmark$ $\checkmark$ - - - - LIFT
CP Others $\checkmark$ - - $\checkmark$ - - CCM
CI + CD Parallel CNN (CI) and BiLSTM/Transformer(CD) based Dual Encoder + ETA $\checkmark$ - - - $\checkmark$ $\checkmark$ MSTN
The established field of multivariate time–series forecasting is conventionally organized around three channel interaction strategies—channel–independent (CI), channel–dependent (CD), and channel–projected (CP) approaches [qiu2025comprehensivesurveydeeplearning]—as summarized in Table 6. These strategies specify how a model encodes and exploits inter-variable dependencies across the $C$ channels of a multivariate signal.
CI. CI methods (e.g., DLinear, PatchTST) process each variable independently using lightweight temporal operators such as linear layers or multilayer perceptrons. Their advantage lies in linear channel cost $\mathcal{O}(C)$ and efficient temporal complexity $\mathcal{O}(L)$, though this comes at the expense of discarding cross-channel structure.
CD. CD approaches (e.g., iTransformer) explicitly model inter-variable correlation through dense attention or graph-based mechanisms. While these models offer comprehensive feature coupling, they incur high computational cost—typically $\mathcal{O}(C^{2})$ or $\mathcal{O}(L^{2})$—which limits their scalability and latency characteristics.
CP. CP models (e.g., MCformer, MTGNN) map the channel dimension into a low-rank latent space before applying interaction, thereby avoiding the full $\mathcal{O}(C^{2})$ cost of CD models. Although channel-efficient, these models often impose simplified temporal assumptions, reducing their ability to capture multi-scale dynamics.
Despite the usefulness of this taxonomy, it does not fully resolve the central challenges in long-sequence time–series applications, where models must simultaneously capture short- and long-range temporal dependencies while ensuring low-latency inference. Consequently, there is a need for architectures capable of representing multi-scale temporal structure without incurring prohibitive computational overhead.
Existing SOTA systems illustrate these trade-offs. CI-based PatchTST [nie2023timeseriesworth64] relies on fixed patch-based tokenization; LLM-derived approaches such as LLM4TS [10.1145/3719207] and TIME-LLM [jin2024timellmtimeseriesforecasting] depend on frozen backbone architectures; CD-based TimesNet [wu2023timesnettemporal2dvariationmodeling] employs fixed frequency priors; and CP-based MCformer [10533212] attains efficiency through partial channel blending. These designs incorporate rigid structural priors that limit their ability to capture diverse temporal scales without sacrificing computational efficiency.
To address these limitations, we introduce MSTN, a hybrid architecture grounded in the Efficient Temporal Aggregation (ETA) principle. MSTN explicitly targets the multi-scale nature of long-sequence data while avoiding the $\mathcal{O}(L^{2})$ bottleneck typical of Transformer-based models, enabling $\mathcal{O}(1)$ inference with respect to sequence length. By combining the complementary strengths of CI and CD models with a novel temporal aggregation mechanism, MSTN delivers strong performance across forecasting, imputation, and classification tasks while maintaining the low-latency characteristics required for real-world deployment.
4.2.1 Core Innovation: Dual-Path Design and ETA
Table 7: Comparison of architectural characteristics. This table contrasts the MSTN framework against the three standard Channel Interaction strategies (CI, CD, CP) across critical performance dimensions. MSTN achieves high capacity while neutralizing the $\mathcal{O}(L^{2})$ computational bottleneck.
Dimension CI CD CP (CI+CD) MSTN Rationale
Efficiency High Low Moderate High ETA enables $\mathcal{O}(1)$ refinement after initial encoding.
Robustness High Low Moderate High Dual-encoder with SGF provides complementary feature stability.
Generaliza- bility Low Moderate High High Achieved via a single, aggregated feature vector ($\mathbf{z}_{\text{final}}$) that supports multi-task learning (Forecasting, Imputation, Classification).
Capacity Low High Moderate High Transformer/BiLSTM path captures global dependencies.
Ease of Implementation High Moderate Low Moderate Relies on standard components (CNN, Transformer/BiLSTM, SGF, SE, MHA) rather than complex, specialized graph or recursive structures.
The MSTN framework integrates a dual-path temporal encoding architecture with the ETA mechanism, yielding a principled hybrid approach that reconciles the core trade-offs among CI-, CD-, and CP-based modelling strategies as shown in Table 7. Its design is centered on two architectural innovations that jointly enhance representational fidelity and computational efficiency:
  1. 1.
    Dual-branch temporal encoding. In contrast to recent CP-oriented architectures such as MCformer [10533212], which prioritise efficiency at the expense of temporal expressiveness, MSTN explicitly preserves multi-scale temporal structure. It deploys a parallel encoding strategy in which a global sequence-modeling pathway (Transformer or BiLSTM; CD strategy) captures long-range temporal dependencies, while a lightweight convolutional pathway (CI strategy; $\mathcal{O}(L)$) extracts fine-grained local temporal patterns. This duality directly addresses the multi-scale nature of real-world temporal signals, offering coverage beyond what is typically afforded by efficient CP-based methods.
  2. 2.
    Efficiency through ETA: collapsing $\mathcal{O}(L^{2})$ to $\mathcal{O}(1)$. MSTN applies the ETA mechanism immediately after the high-capacity encoders, collapsing the temporal dimension ($L\to 1$) via learned sequence aggregation. By performing the computationally intensive operations (e.g., $\mathcal{O}(L^{2})$ Transformer self-attention) before this aggregation step, the subsequent refinement layers—including feature fusion, squeeze–excitation recalibration, multi-head attention, and prediction modules—operate with a fixed $\mathcal{O}(1)$ cost with respect to $L$. This explicitly reshapes the model’s complexity profile, enabling MSTN to maintain the representational power of CD architectures while achieving substantially improved inference efficiency.
4.3 Baselines
This section presents the baselines used to assess MSTN on four time series tasks: (1) imputation, (2) long-term forecasting, (3) classification, and (4) generalizability study.
4.3.1 Imputation
Following the TimesNet [wu2023timesnettemporal2dvariationmodeling] protocol, we evaluate performance under the MCAR scenario with random masking ratios of $\{12.5\%,25\%,37.5\%,50\%\}$. Performance is quantified using MSE and MAE. The evaluation includes recent SOTA baselines such as GPT2(3) [zhou2023fitsallpowergeneraltime], TimesNet [wu2023timesnettemporal2dvariationmodeling], PatchTST [nie2023timeseriesworth64], LightTS [zhang2022morefastmultivariatetime], and DLinear [zeng2022transformerseffectivetimeseries].
4.3.2 Long Term Forecasting
The proposed MSTN model is evaluated for the long-term time series forecasting task. We evaluated all methods across multiple prediction horizons $H\in\{96,192,336,720\}$ for standard benchmarks and $H\in\{24,36,48,60\}$ for the ILI dataset. The forecast performance is quantified using MSE and MAE. The evaluation includes the following SOTA baselines: EMTSF [alharthi2025emtsfextraordinarymixturesotamodels], LLM4TS [10.1145/3719207], HiMTM [zhao2024himtmhierarchicalmultiscalemasked], TIME-LLM [jin2024timellmtimeseriesforecasting], MTST [zhang2024multiresolutiontimeseriestransformerlongterm], SOFTS [han2024softsefficientmultivariatetime], and iTransformer [liu2024itransformerinvertedtransformerseffective].
4.3.3 Classification
The performance of classification tasks is evaluated using the standard metric of classification accuracy. The evaluation compares with recent SOTA baselines such as GPT2(6) [alharthi2025emtsfextraordinarymixturesotamodels], TimesNet [wu2023timesnettemporal2dvariationmodeling], DLinear [zeng2022transformerseffectivetimeseries], ETSFormer [woo2022etsformer], FEDformer [zhou2022fedformer], and Informer [zhou2021informer].
4.3.4 Generalizability Study
To evaluate generalizability between domains of the proposed MSTN framework, we benchmarked against baselines (e.g. TimesNet [wu2023timesnettemporal2dvariationmodeling], PatchTST [nie2023timeseriesworth64]) on seven publicly available international datasets spanning healthcare, activity recognition, agricultural technology, and industrial monitoring. All evaluations utilized a consistent five-fold cross-validation framework to ensure a robust and fair comparison.
4.4 Experimental Setup
4.4.1 Training and Model Configuration
The proposed MSTN model is implemented in Python 3.13.1 using PyTorch 2.7.1 and trained on an NVIDIA T400 GPU (4 GB VRAM). The architecture processes variable length sequences through the sequence modeling cores BiLSTM (128 hidden units) or Transformer (4 layers, 8 attention heads), combined with a CNN pathway (128 $\to$ 64 filters). Training uses AdamW [kingma2017adammethodstochasticoptimization] optimizer with learning rate $3\times 10^{-4}$, batch size 64, and task-specific objectives (Focal Loss for classification, MSE for forecasting and imputation). Models train for up to 100 epochs with early stopping based on validation performance. Data preprocessing employs min-max normalization to standardize feature distributions. For forecasting and imputation benchmarks, we utilize a fixed input lookback window of $L=96$ (except $L=36$ for ILI) across all prediction horizons, whereas for classification tasks involving variable length sequences (e.g., JapaneseVowels), inputs are standardized to the maximum sequence length (e.g., $L=29$) to facilitate consistent batch processing.
4.4.2 Evaluation Metrics
The performance of forecasting and imputation tasks is evaluated using two primary regression metrics that measure the discrepancy between predicted values and ground truth. For forecasting tasks, these metrics quantify the accuracy of future predictions, while for imputation tasks they assess the quality of missing value reconstruction. The performance is evaluated using:
Mean Squared Error: $\quad\text{MSE}=\frac{1}{H}\sum_{i=1}^{H}(Y_{i}-\hat{Y}_{i})^{2}$ (15)
Mean Absolute Error: $\quad\text{MAE}=\frac{1}{H}\sum_{i=1}^{H}|Y_{i}-\hat{Y}_{i}|$ (16)
where $Y,\hat{Y}\in\mathbb{R}^{H\times C}$ are the ground truth and prediction results of the future with $H$ time points and $C$ channels. $Y_{i}$ denotes the $i$-th future time point. The performance of classification tasks is evaluated using four standard classification metrics that measure different aspects of predictive performance, particularly important for handling potential class imbalance in the datasets. The classification performance is evaluated using:
$\text{Accuracy}=\frac{\text{TP}+\text{TN}}{\text{TP}+\text{TN}+\text{FP}+\text{FN}}$ (17)
$\text{Precision}=\frac{\text{TP}}{\text{TP}+\text{FP}}$ (18)
$\text{Recall}=\frac{\text{TP}}{\text{TP}+\text{FN}}$ (19)
$\text{F1-Score}=2\times\frac{\text{Precision}\times\text{Recall}}{\text{Precision}+\text{Recall}}$ (20)
where TP denotes True Positives, TN True Negatives, FP False Positives, and FN False Negatives.
5 Acknowledgments
This research is supported by the IIT Indore Young Faculty Research Catalysing Grant (YFRCG) Scheme ( Project ID: IITI/YFRCG/2023-24/01), Indian Institute of Technology Indore, India.
6 Data Availability
All datasets used in this study are publicly available from their respective sources as cited throughout the paper.
7 Author Contributions
Sumit S Shevtekar led the research design, developed the architecture, implemented all software, conducted experiments, performed data analysis and visualization, and wrote and edited the original draft. Chandresh K Maurya contributed to experimental validation, manuscript writing, review, and editing